# Sharp concentration inequality for the sum of random variables

### v1

@@
Abstract
We present a universal concentration bound for sums of random variables under arbitrary dependence, and we prove it is asymptotically optimal for every fixed common marginal law. The concentration bound is a direct—yet previously unnoticed—consequence of the subadditivity of expected shortfall, a property well known to financial statisticians. The sharpness result is a significant contribution relying on the construction of worst-case dependency profiles between identically distributed random variables.
1 Maximally nonincreasing operator framework and main results
Following the approach of Rockafellar and Royset (2014) on superquantiles and our previous work on concentration inequalities Louart (2024), we encode concentration statements as inequalities between maximally nonincreasing operators rather than as pointwise inequalities between real-valued functions. The operator viewpoint is convenient here because objects underlying tail bounds (like survival functions and quantiles) are monotone but may have jumps (like the limiting profile of the law of large numbers); representing them as set-valued operators provides a canonical way to handle discontinuities without choosing arbitrary versions (lower or upper semicontinuous).
Concretely, we work in the class of “maximally nonincreasing operators” generalizing nonincreasing mappings on $\mathbb{R}$, that we denote $\mathcal{M}_{\downarrow}$ and that contains set-valued mappings $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ satisfying [^1] The domain, range, and graph of an operator $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ are respectively defined as $\operatorname{Dom}(\alpha)=\{t\in\mathbb{R}\ |\ \alpha(t)\neq\emptyset\}$, $\operatorname{Ran}(\alpha)=\cup_{t\in\mathbb{R}}\alpha(t)$, and $\operatorname{Gra}(\alpha)=\{(x,y)\,:\,x\in\operatorname{Dom}(\alpha),\ y\in\alpha(x)\}$. The domain and range of a maximally nonincreasing operator are intervals; moreover, for any $t\in\mathbb{R}$, $\alpha(t)$ is a closed interval.:
@@ 1 Maximally nonincreasing operator framework and main results
Further characterizations can be found in Bauschke and Combettes (2011); Louart (2024).
In what follows, $\mu$ designates a probability measure on $\mathbb{R}$ and $X\sim\mu$, a given random variable. Operator inversion allows one to move naturally between the survival operator $S_{X}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $S_{\mu}$) and the tail quantile operator $T_{X}:=S_{X}^{-1}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $T_{\mu}$), both maximally nonincreasing and defined respectively by
$\forall t\in\mathbb{R}:\qquad$ $S_{X}(t):=[\mathbb{P}(X>t),\mathbb{P}(X\geq t)]=[\mu(]t,+\infty[),\mu(]t,+\infty[)]=:S_{\mu}(t)$
@@
The natural order relation to set concentration inequalities on survival operator was introduced in Louart (2024):
Definition 1 (Interval Order and Point-wise resolvent Order between operators).
The order between intervals $A,B\subset\mathbb{R}$ is defined by
@@
where $\operatorname{Dom}(f)$, $\operatorname{Dom}(g)$, $f(x)$ and $g(x)$ are (by construction) intervals (see (Bauschke and Combettes, 2011, Proposition 20.31, Corollary 21.12)).
A confortable advantage of working with survival and tail quantile operators instead of working with cumulative distribution operator $t\mapsto 1-S_{X}(t)$ and quantile operator $p\mapsto T_{X}(1-p)$ as in Rockafellar and Royset (2014) is that an upper bound on the survival profile is equivalent to the corresponding upper bound on the tail quantile profile. More precisely, for any $\alpha\in\mathcal{M}_{\downarrow}$,
@@
Our major task then expresses:
@@
    Given $n$ probability measures $\mu_{1},\ldots,\mu_{n}$ on $\mathbb{R}$, construct a minimal operator $\beta_{\mu_{1},\ldots,\mu_{n}}\in\mathcal{M}_{\downarrow}$ such that, for any collection of $n$ random variables $X_{1}\sim\mu_{1},\ldots,X_{n}\sim\mu_{n}$:
    $S_{X_{1}+\cdots+X_{n}}\leq\beta_{\mu_{1},\ldots,\mu_{n}}.$
Assuming $\mu_{1}=\cdots=\mu_{n}$ and (i) considering $X_{1},\ldots,X_{n}$ i.i.d, (ii) considering $X_{1}=\cdots=X_{n}$ and (iii) using the union bound, on gets the first simple bound [^2] Maximum of operators is given a rigorous definition in Louart (2024) that is not necessary to provide here since this inequality just plays a heuristic role.
$\max(\mathrm{Incr}_{\mathbb{E}[X]},S_{X})\leq\beta_{\mu,\ldots,\mu}\leq nS_{X},$
where $X\sim\mathbb{\mu}$ and for any $\delta\in\mathbb{R}$, we denote by $\mathrm{Incr}_{\delta}\in\mathcal{M}_{\downarrow}$ the operator defined by
@@
  • denoting $p_{f}:=\sup(\operatorname{Dom}(f))$, we have $p_{f}\in\operatorname{Dom}(\mathcal{H}(f))$ and $\mathcal{H}(f)(p_{f})=\left]-\infty,\frac{1}{p_{f}}\int_{0}^{p_{f}}f(r)\,dr\right]$.
An example of a Hardy transform is depicted on Figure 1. We record a few properties that will be useful later:
[…]
    $0,1]\subset\operatorname{Dom}(\mathcal{H}(f))$ and for all $t\in]0,1[$, $\mathcal{H}(f)(t)$ is a singleton;
[…]
@@
A first partial answer to our Task is then given by
Theorem 1.
@@
As a comparison, in (Louart, 2024, Proposition 42) we only proved the weaker bound (a generalization of union bound to the non-identically distributed case):
@@
When $X_{1},\ldots,X_{n}\sim\mu$ are identically distributed, Theorem 1 rewrites:
@@
Remark 1.
It is instructive to understand what is lost at the level of moments when applying Theorem 1. For any nonnegative random variable $X\sim\mu$, for some probability measure $\mu$ on $\mathbb{R}$ one has the identity (see for instance (Rockafellar and Royset, 2014, (3.4))):
$M_{q}(\mu):=\mathbb{E}[X^{q}]=\int_{0}^{\infty}\mathbb{P}(X^{q}\geq t)\,dt=\int_{0}^{\infty}T_{X}(p)^{q}\,dp.$ (3)
Following Hardy’s inequality (Hardy et al., 1952, Theorem 9.8.2), valid for any measurable function $f:]0,1[\to\mathbb{R}_{+}$,
@@
Given an operator $\beta\in\mathcal{M}_{\downarrow}$ with $\operatorname{Dom}(\beta)\subset\mathbb{R}_{+}$, we define the “moments” of $\beta$ by
$M_{q}(\beta):=\int_{\mathbb{R}_{+}}\beta\!\left(t^{\frac{1}{q}}\right)\,dt,\quad q>0,$ (4)
so that $M_{q}(S_{X})=\mathbb{E}[X^{q}]=M_{q}(\mu)$. Then, given $n$ random variables $X_{1},\ldots,X_{n}\sim\mu$, Theorem 1 applied in the $(X_{i})_{i\in[k]}$-identically distributed setting (2) yields the moment bound
$M_{q}\left(S_{\frac{1}{n}\sum_{i=1}^{n}X_{k}}\right)$ $=\int_{0}^{1}T_{\frac{1}{n}\sum_{i=1}^{n}X_{k}}(p)^{q}\,dp\leq\int_{0}^{1}\mathcal{H}(T_{X})(p)^{q}\,dp$
$\leq\left(\frac{q}{q-1}\right)^{q}\int_{0}^{1}T_{X}(p)^{q}\,dp=\left(\frac{q}{q-1}\right)^{q}M_{q}(S_{X})$ (5)
This inequality is clearly underperforming since a mere application of Jensen inequality already yields for all $q>1$ the moment optimal bound [^6] It is reached for instance in the case $X_{1}=\cdots=X_{n}$.:
$M_{q}(S_{\frac{1}{n}\sum_{i=1}^{n}X_{k}})\leq M_{q}(S_{X})$
Still (1) demonstrates that very little is lost through application of Theorem 1. Anyway, the strength of Theorem 1 is that it provides a universal tail bound, valid for all thresholds $t$ and all joint distributions of $(X_{1},\ldots,X_{n})$ with given marginals; for this reason, moments are not the most relevant metric here.
To pursue the study of the scope of Theorem 1, we give below a general result that efficiently bounds the concentration of the sum for a wide range of distributions. We employ the shorthand notation (already introduced in Louart (2024)) $\mathrm{Id}^{-a}:\mathbb{R}\to 2^{\mathbb{R}}$ defined by
$\forall t>0:\quad\mathrm{Id}^{-a}(t)=\{t^{-a}\},\qquad\text{and}\qquad\mathrm{Id}^{-a}((-\infty,0])=\emptyset.$
Corollary 1.
Given a probability measure $\mu$ on $\mathbb{R}$ and $n$ identically distributed random variables $X_{1},\ldots,X_{n}\sim\mu$, assume that there exist $\alpha\in\mathcal{M}_{\downarrow}$ satisfying $\operatorname{Ran}(\alpha)\subset\mathbb{R}^{+}_{*}$ and $q>1$ such that [^7] We say that an operator $g:\mathbb{R}\to 2^{\mathbb{R}}$ is convex if any real-valued mapping $h:\operatorname{Dom}(g)\to\mathbb{R}$ satisfying $\forall t\in\operatorname{Dom}(g)$, $h(t)\in g(t)$ is convex.:
$S_{\mu}\leq\alpha\qquad\text{and}\qquad\mathrm{Id}^{-\frac{1}{q}}\circ\alpha\ \ \text{is convex}.$
Then one can bound
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq\left(\frac{q}{q-1}\right)^{q}\alpha.$
If we assume in addition [^8] The second result of Theorem 3 in the next section states that $\log\circ\alpha$ convex implies $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ convex for all $q>0$. that $\log\circ\alpha$ is convex, then
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq e\,\alpha.$
@@
Theorem 1 is, in fact, a reformulation of a property that is classical in financial mathematics: the subadditivity of expected shortfall Acerbi and Tasche (2002) (also called conditional value-at-risk Rockafellar and Uryasev (2002), and more recently superquantiles Rockafellar and Royset (2014)). Subadditivity is one of the coherence axioms introduced in Artzner et al. (1999) for risk measures. This property was obtained simultaneously in (Acerbi and Tasche, 2002, Proposition 3.1) (under the name “shortfall expectation”) and in (Rockafellar and Uryasev, 2002, Corollary 12) (under the name “conditional value-at-risk”). Although these notions were initially presented differently, they were later shown to coincide; in our notation, they correspond to the mapping $p\mapsto\mathcal{H}(T_{X})(1-p)$. Subadditivity can be written as
Noting that $T_{X_{1}+\cdots+X_{n}}\leq\mathcal{H}(T_{X_{1}+\cdots+X_{n}})$ and relying on (1), we recover exactly Theorem 1.
Up to this point, no structurally new phenomenon has been introduced; we now have all the ingredients to state our main contribution: asymptotic sharpness of the inequality provided by Theorem 1. Before stating the result, if $\mathbb{E}[|X|]<\infty$, let us introduce the quantity
@@
Proof.
First note that $\Delta_{\mu}(p)\geq 0$ since $\mathcal{H}(T_{\mu})$ is nonincreasing and $\mathcal{H}(T_{\mu})(p)\geq\mathcal{H}(T_{\mu})(1)=\mathbb{E}[X]$. Moreover,
@@
which proves (7) and the characterization of the equality case. ∎
Define the limiting survival operator $S_{\mu,p}\in\mathcal{M}_{\downarrow}$ satisfying on continuity points:
$S_{\mu,p}(t)=\begin{cases}\{1\},&\text{if }t<H(T_{\mu})(p)-\Delta_{\mu}(p),\\ \{p\},&\text{if }H(T_{\mu})(p)-\Delta_{\mu}(p)<t<H(T_{\mu})(p),\\ \{0\},&\text{if }t>H(T_{\mu})(p).\end{cases}$
Finally, we set $S_{\mu,1}:=\mathrm{Incr}_{\mathbb{E}[X]}$ (note that $\mathcal{H}(T_{\mu})(1)=\mathbb{E}[X]$). In particular, the sharpness will be shown at the contact point:
$p\in S_{\mu,p}(H(T_{\mu})(p)),$
that we represented by a green dot on the tail-quantile profiles graph provided on Figure 1
@@
A final answer to our Task (in the identically distributed and finite expectation case) is given by
Theorem 2.
Given a probability distribution $\mu$ on $\mathbb{R}$ admitting a finite expectation, and given any $p\in(0,1]$, there exists a sequence of identically distributed random variables $(X_{n})_{n\in\mathbb{N}}\in\mathbb{R}^{\mathbb{N}}$ such that for all $n\in\mathbb{N}$, $X_{n}\sim\mu$ and [^9] That is, for any continuity point of $S_{\mu,p}$. for all $t\in\mathbb{R}\setminus\{a_{\mu}(p),b_{\mu}(p)\}$,
$S_{\frac{1}{n}(X_{1}+\cdots+X_{n})}(t)\underset{n\to\infty}{\longrightarrow}S_{\mu,p}(t).$
This theorem will be proved constructively in Section 2.

### v2

# Let $X_{1},\ldots,X_{n}$ be random variables with an arbitrary dependence structure. Assume that there exists a nonincreasing mapping $\alpha:\mathbb{R}\to[0,1]$ with $\alpha(\mathbb{R})=[0,1]$ such that

@@
[…]
1 Introduction and main results
Let $X_{1},\ldots,X_{n}$ be random variables with an arbitrary dependence structure. Assume that there exists a nonincreasing mapping $\alpha:\mathbb{R}\to[0,1]$ with $\alpha(\mathbb{R})=[0,1]$ such that
$\forall i\in[n],\ \forall t\in\mathbb{R}:\qquad\mathbb{P}(X_{i}\geq t)\leq\alpha(t).$
Then, as a direct consequence of a well-known property in the risk-measure literature, namely the subadditivity of expected shortfall, one obtains the universal bound (see Theorem 1.2)
$\mathbb{P}\!\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq t\right)\leq\bigl(\mathcal{H}(\alpha^{-1})\bigr)^{-1}(t),$
where $\alpha^{-1}$ denotes the generalized inverse of $\alpha$ and $\mathcal{H}(\alpha^{-1})$ is the Hardy transform, defined by
$\forall u\in(0,1]:\qquad\mathcal{H}(\alpha^{-1})(u):=\frac{1}{u}\int_{0}^{u}\alpha^{-1}(p)\,dp.$
The main contribution of this paper is to show, by an explicit construction, that without any additional assumption on the dependence structure the bounding profile $\bigl(\mathcal{H}(\alpha^{-1})\bigr)^{-1}$ is asymptotically optimal as $n\to\infty$ when $\alpha$ is the survival profile of the marginal law (see Theorems 1.82.1). More broadly, the operator formalism developed below covers nonincreasing, possibly discontinuous, survival profiles $\alpha$, and it naturally accommodates heterogeneous marginals, yielding a sharp universal analogue of the union bound for sums.
Following the approach of [Rockafellar and Royset(2014)] on superquantiles and our previous work on concentration inequalities [Louart(2024)], we encode concentration statements as inequalities between maximally nonincreasing operators rather than as pointwise inequalities between real-valued functions. The operator viewpoint is convenient here because objects underlying tail bounds (like survival functions and quantiles) are monotone but may have jumps (like the limiting profile of the law of large numbers); representing them as set-valued operators provides a canonical way to handle discontinuities without choosing arbitrary versions (lower or upper semicontinuous). In particular, atoms force a choice of left/right versions for survival functions and quantiles, and these choices can obscure inversion identities; interval-valued operators avoid this issue.
Concretely, we work in the class of “maximally nonincreasing operators” generalizing nonincreasing mappings on $\mathbb{R}$, which we denote $\mathcal{M}_{\downarrow}$ and which consists of set-valued mappings $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ satisfying [^2] The domain, range, and graph of an operator $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ are respectively defined as $\operatorname{Dom}(\alpha)=\{t\in\mathbb{R}\ |\ \alpha(t)\neq\emptyset\}$, $\operatorname{Ran}(\alpha)=\cup_{t\in\mathbb{R}}\alpha(t)$, and $\operatorname{Gra}(\alpha)=\{(x,y)\,:\,x\in\operatorname{Dom}(\alpha),\ y\in\alpha(x)\}$. The domain and range of a maximally nonincreasing operator are intervals; moreover, for any $t\in\mathbb{R}$, $\alpha(t)$ is a closed interval.:
@@ 1 Introduction and main results
Further characterizations can be found in [Bauschke and Combettes(2011), Louart(2024)].
In what follows, $\mu$ designates a probability measure on $\mathbb{R}$ and $X\sim\mu$ a generic random variable. Operator inversion allows one to move naturally between the survival operator $S_{X}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $S_{\mu}$) and the tail quantile operator $T_{X}:=S_{X}^{-1}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $T_{\mu}$), both maximally nonincreasing and defined respectively by
$\forall t\in\mathbb{R}:\qquad$ $S_{X}(t):=[\mathbb{P}(X>t),\mathbb{P}(X\geq t)]=[\mu(]t,+\infty[),\,\mu([t,+\infty[)]=:S_{\mu}(t)$
@@
In particular, at an atom $t_{0}$, $S_{X}(t_{0})$ is the interval $\mathbb{P}(X>t_{0}),\mathbb{P}(X\geq t_{0})]$.
The natural order relation used to state concentration inequalities for survival operators was introduced in [Louart(2024)]:
Definition 1.1 (Interval Order and Point-wise resolvent Order between operators).
The order between intervals $A,B\subset\mathbb{R}$ is defined by [^3] For instance, if $A=[a_{1},a_{2}]$ and $B=[b_{1},b_{2}]$, then: $A\leq B\Leftrightarrow a_{1}\leq b_{1}\ \text{and}\ a_{2}\leq b_{2}$.
@@
where $\operatorname{Dom}(f)$, $\operatorname{Dom}(g)$, $f(x)$ and $g(x)$ are (by construction) intervals (see [Bauschke and Combettes(2011), Proposition 20.31, Corollary 21.12]).
A convenient advantage of working with survival and tail quantile operators instead of working with the cumulative distribution operator $t\mapsto 1-S_{X}(t)$ and the quantile operator $p\mapsto T_{X}(1-p)$ as in [Rockafellar and Royset(2014)] is that an upper bound on the survival profile is equivalent to the corresponding upper bound on the tail quantile profile. More precisely, for any $\alpha\in\mathcal{M}_{\downarrow}$,
@@
Our objective can be stated as follows:
@@
    Given $n$ probability measures $\mu_{1},\ldots,\mu_{n}$ on $\mathbb{R}$, construct a minimal [^4] For the pointwise resolvent order defined in Definition 1.1. operator $\alpha_{\mu_{1},\ldots,\mu_{n}}\in\mathcal{M}_{\downarrow}$ such that, for any collection of $n$ random variables $X_{1}\sim\mu_{1},\ldots,X_{n}\sim\mu_{n}$:
    $S_{X_{1}+\cdots+X_{n}}\leq\alpha_{\mu_{1},\ldots,\mu_{n}}.$
Assuming $\mu_{1}=\cdots=\mu_{n}=\mu$ and (i) considering $X_{1},\ldots,X_{n}$ i.i.d., (ii) considering $X_{1}=\cdots=X_{n}$ and (iii) using the union bound, one obtains the first simple bound [^5] Maximum of operators is given a rigorous definition in [Louart(2024)] that is not necessary to provide here since this inequality just plays a heuristic role.
$\max(\mathrm{Incr}_{\mathbb{E}[X]},S_{X})\leq\alpha_{\mu,\ldots,\mu}\leq nS_{X},$
where for any $\delta\in\mathbb{R}$, we denote by $\mathrm{Incr}_{\delta}\in\mathcal{M}_{\downarrow}$ the operator defined by
@@
The naive upper bound $nS_{X}$ can of course be substantially improved. The optimal bound is expressed through the Hardy transform of the tail quantile operator $T_{X}$.
@@
  • denoting $p_{f}:=\sup(\operatorname{Dom}(f))$, we have $p_{f}\in\operatorname{Dom}(\mathcal{H}(f))$ and:
    $\mathcal{H}(f)(p_{f})=\left]-\infty,\frac{1}{p_{f}}\int_{0}^{p_{f}}f(r)\,dr\right]$
An example of a Hardy transform is depicted on Figure 1. We record a few elementary properties that will be useful later:
  • $0,p_{f}]\subset\operatorname{Dom}(\mathcal{H}(f))$ and for all $t\in]0,p_{f}[$, $\mathcal{H}(f)(t)$ is a singleton.
@@
A first partial answer to our Task is, in fact, a reformulation of a property that is classical in financial mathematics: the subadditivity of expected shortfall [Acerbi and Tasche(2002)] (also called conditional value-at-risk [Rockafellar and Uryasev(2002)], and more recently superquantiles [Rockafellar and Royset(2014)]). Subadditivity is one of the coherence axioms introduced in [Artzner et al.(1999)Artzner, Delbaen, Eber, and Heath] for risk measures. This property was obtained simultaneously in [Acerbi and Tasche(2002), Proposition 3.1] (under the name “shortfall expectation”) and in [Rockafellar and Uryasev(2002), Corollary 12] (under the name “conditional value-at-risk”). Although these notions were initially presented differently, they were later shown to coincide; in our notation, they correspond to the mapping $p\mapsto\mathcal{H}(T_{X})(1-p)$. Subadditivity can be written as
Noting that $T_{X_{1}+\cdots+X_{n}}\leq\mathcal{H}(T_{X_{1}+\cdots+X_{n}})$ and relying on (1), we recover exactly:
Theorem 1.2.
@@
As a comparison, in [Louart(2024), Proposition 42] we only set the weaker elementary bound (a generalization of the union bound to the non-identically distributed case):
@@
When $X_{1},\ldots,X_{n}\sim\mu$ are identically distributed, Theorem 1.2 rewrites:
@@
Example 1.3 (Two-point distribution).
Let $X\sim\mu$ take only two values $a<b$, with
$\mathbb{P}(X=b)=\pi,\qquad\mathbb{P}(X=a)=1-\pi,\qquad\pi\in(0,1),$
so $\mathbb{E}[X]=a+\pi(b-a)$. Applying Theorem 1.2 yields the survival bound
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}(t)\leq\mathcal{H}(T_{\mu})^{-1}(t)=\begin{cases}\{1\},&t<\mathbb{E}[X],\\[4.0pt] \displaystyle\left\{\frac{\pi(b-a)}{t-a}\right\},&\mathbb{E}[X]<t<b,\\[8.0pt] [0,\pi],&t=b,\\[2.0pt] \{0\},&t>b.\end{cases}$
Remark 1.4.
While Theorem 1.2 is not designed to be tight at the level of moments, it is instructive to quantify what is lost when translating a tail bound into a moment bound. Still, it must be kept in mind that the strength of Theorem 1.2 is that it provides a universal tail bound, valid for all thresholds $t$ and all joint distributions of $(X_{1},\ldots,X_{n})$ with given marginals; for this reason, moments are not the most relevant metric here.
For any nonnegative random variable $X\sim\mu$, for some probability measure $\mu$ on $\mathbb{R}$ one has the identity (see for instance [Rockafellar and Royset(2014), (3.4)]):
$M_{q}(\mu):=\mathbb{E}[X^{q}]=\int_{0}^{\infty}\mathbb{P}(X^{q}\geq t)\,dt=\int_{0}^{1}T_{X}(p)^{q}\,dp.$ (3)
Following Hardy’s inequality [Hardy et al.(1952)Hardy, Littlewood, and Pólya, Theorem 9.8.2], valid for any measurable function $f:]0,1[\to\mathbb{R}_{+}$,
@@
Given an operator $\alpha\in\mathcal{M}_{\downarrow}$ with $\operatorname{Dom}(\alpha)\subset\mathbb{R}_{+}$, we define the “moments” of $\alpha$ by
$M_{q}(\alpha):=\int_{\mathbb{R}_{+}}\alpha\!\left(t^{\frac{1}{q}}\right)\,dt,\quad q>0,$
so that $M_{q}(S_{X})=\mathbb{E}[X^{q}]=M_{q}(\mu)$.
Then, given $n$ random variables $X_{1},\ldots,X_{n}\sim\mu$, Theorem 1.2 applied in the $(X_{i})_{i\in[n]}$-identically distributed setting (2) yields the moment bound
$M_{q}\left(S_{\frac{1}{n}\sum_{i=1}^{n}X_{i}}\right)$ $=\int_{0}^{1}T_{\frac{1}{n}\sum_{i=1}^{n}X_{i}}(p)^{q}\,dp\leq\int_{0}^{1}\mathcal{H}(T_{X})(p)^{q}\,dp$
$\leq\left(\frac{q}{q-1}\right)^{q}\int_{0}^{1}T_{X}(p)^{q}\,dp=\left(\frac{q}{q-1}\right)^{q}M_{q}(S_{X}).$ (4)
This inequality is not tight for moments since a mere application of Jensen’s inequality already yields for all $q>1$ the moment optimal bound [^9] It is reached for instance in the case $X_{1}=\cdots=X_{n}$.:
$M_{q}(S_{\frac{1}{n}\sum_{i=1}^{n}X_{i}})\leq M_{q}(S_{X}).$
To pursue the study of the scope of Theorem 1.2, we give below a general result that efficiently bounds the concentration of the sum for a wide range of distributions. We employ the shorthand notation (already introduced in [Louart(2024)]) $\mathrm{Id}^{-a}:\mathbb{R}\to 2^{\mathbb{R}}$ defined by
$\forall t>0:\quad\mathrm{Id}^{-a}(t)=\{t^{-a}\},\qquad\text{and}\qquad\mathrm{Id}^{-a}((-\infty,0])=\emptyset.$ (5)
Corollary 1.5.
Given $n$ random variables $X_{1},\ldots,X_{n}$ and $\alpha\in\mathcal{M}_{\downarrow}$ satisfying $\operatorname{Ran}(\alpha)\subset\mathbb{R}^{+}_{*}$ and $q>1$ such that [^10] We say that an operator $g:\mathbb{R}\to 2^{\mathbb{R}}$ is convex if any real-valued mapping $h:\operatorname{Dom}(g)\to\mathbb{R}$ satisfying $\forall t\in\operatorname{Dom}(g)$, $h(t)\in g(t)$ is convex.:
$\forall i\in[n],\forall t\in\mathbb{R}:\quad\mathbb{P}(X_{i}\geq t)\leq\alpha(t)\qquad\text{and}\qquad\mathrm{Id}^{-\frac{1}{q}}\circ\alpha\ \ \text{is convex},$
one can bound
$\forall t\in\mathbb{R}:\quad\mathbb{P}\left(\frac{1}{n}\sum_{k=1}^{n}X_{k}\geq t\right)\leq\left(\frac{q}{q-1}\right)^{q}\alpha(t).$
If we assume in addition [^11] The second result of Proposition 3.2 in the next section states that $\log\circ\alpha$ convex implies $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ convex for all $q>0$. that $\log\circ\alpha$ is convex, then
$\forall t\in\mathbb{R}:\quad\mathbb{P}\left(\frac{1}{n}\sum_{k=1}^{n}X_{k}\geq t\right)\leq e\alpha(t).$
@@
We now turn to our main contribution: asymptotic sharpness of the inequality provided by Theorem 1.2. A general result for a triangular array of marginals $(\mu_{1}^{(n)},\ldots,\mu_{n}^{(n)})$ is stated in Section 2 (see Theorem 2.1); to keep the picture as simple as possible in this introduction, we first focus on the identically distributed case, where for each $n\in\mathbb{N}$ one has $\mu_{1}^{(n)}=\cdots=\mu_{n}^{(n)}=\mu$.
Before stating the result (under the assumption $\mathbb{E}[|X|]<\infty$), let us introduce the quantity
@@
Proof 1.7.
First note that $\Delta_{\mu}(p)\geq 0$ since $\mathcal{H}(T_{\mu})$ is nonincreasing and $\mathcal{H}(T_{\mu})(p)\geq\mathbb{E}[X]$ for $p\in(0,1)$. Moreover,
@@
which proves (7) and the characterization of the equality case.
Given $p\in(0,1)$, define the limiting survival operator [^12] If $p=1$, $a_{\mu}(p)=b_{\mu}(p)=\mathbb{E}[X]$ and one can simply define $S_{\mu,1}:=\mathrm{Incr}_{\mathbb{E}[X]}$. $S_{\mu,p}\in\mathcal{M}_{\downarrow}$:
$S_{\mu,p}(t)=\begin{cases}\{1\},&\text{if }t<a_{\mu}(p),\\ [p,1],&\text{if }t=a_{\mu}(p),\\ \{p\},&\text{if }a_{\mu}(p)<t<b_{\mu}(p),\\ [0,p],&\text{if }t=b_{\mu}(p),\\ \{0\},&\text{if }t>b_{\mu}(p),\end{cases}$ $\text{with:}\quad\left\{\begin{aligned} a_{\mu}(p)&:=\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p),\\ b_{\mu}(p)&:=\mathcal{H}(T_{\mu})(p).\end{aligned}\right$.
On the tail-quantile plot in Figure 1, the graph of $S_{\mu,p}^{-1}$ meets the graph of $\mathcal{H}(T_{\mu})$ at the point with abscissa $p$, shown as a green dot. The next theorem shows that such contact points can be achieved for any $p\in(0,1)$, so that the bound of Theorem 1.2 is asymptotically tight in this sense.
@@
Theorem 1.8.
Given a probability distribution $\mu$ on $\mathbb{R}$ admitting a finite expectation, and given any $p\in(0,1]$, there exists for each $n\in\mathbb{N}$ a family of identically distributed random variables
$(X^{(n)}_{1},\ldots,X^{(n)}_{n})\qquad\text{with}\qquad X^{(n)}_{k}\sim\mu\ \text{for all }k\in[n],$
such that [^13] The set $\mathbb{R}\setminus\{a_{\mu}(p),b_{\mu}(p)\}$ is exactly the set of continuity points of $S_{\mu,p}$. for all $t\in\mathbb{R}\setminus\{a_{\mu}(p),b_{\mu}(p)\}$,
$S_{\frac{1}{n}\sum_{k=1}^{n}X^{(n)}_{k}}(t)\underset{n\to\infty}{\longrightarrow}S_{\mu,p}(t).$
This theorem is a direct consequence of Theorem 2.1 stated in Section 2.